which is the solution set for the equation : |x + 4| + 4 = 10

Answer by RAY100(1637)

[x+4] +4 =10
[x+4] =6
.
(x+4) = +6,,,,,,,,,,,(x+4)=-6
x=2,,,,,,,,,,,,,,,,,,,x=-10
.
check,[2+4] +4 = 10,,ok
[ -10 +4] +4 =10,,,[-6] +4 =10,,,6+4 =10,,,,,ok